find the coordinates of point d on line bc such that ad is the shortest segment distance from a to bc for a(9,9) b(-1,11) c(-2,2)
Answer by ewatrrr(24785)


Hi
line containing points (-1,11) and (-2,2). m = 9/1
y = 9x + b |using pt(-2,2) to solve for b
2 = -18 + b
20 = b
y = 9x + 20
9 = 9x + 20 | substituting y =9 to find x value
-11/9 = x
coordinates of point d are (-11/9,9)